Cullumber and Paul Company
Income Statement
For the Year Ended December 31, 2020

Sales (270,000 units) $1,350,000
Cost of goods sold 900,000
Gross profit 450,000
Operating expenses
Selling $315,000
Administrative 175,500
490,500
Net loss $(40,500 )


A cost behavior analysis indicates that 75% of the cost of goods sold are variable, 42% of the selling expenses are variable, and 40% of the administrative expenses are variable.

Cullumber has proposed a plan to get the partnership “out of the red” and improve its profitability. She feels that the quality of the product could be substantially improved by spending $0.25 more per unit on better raw materials. The selling price per unit could be increased to only $5.25 because of competitive pressures. Cullumber estimates that sales volume will increase by 25%. Compute the net income under Cullumber's proposal and the break-even point in dollars. (Round intermediate calculations to 4 decimal places, e.g. 15.2515 and final answers to 0 decimal places, e.g. 2,520.)

Amount
Net income $

Entry field with incorrect answer

Break-even point $

Entry field with incorrect answer now contains modified data

Paul was a marketing major in college. He believes that sales volume can be increased only by intensive advertising and promotional campaigns. He therefore proposed the following plan as an alternative to Cullumber’s: (1) increase variable selling expenses to $0.59 per unit, (2) lower the selling price per unit by $0.25, and (3) increase fixed selling expenses by $43,000. Paul quoted an old marketing research report that said that sales volume would increase by 60% if these changes were made. Compute the net income under Paul’s proposal and the break-even point in dollars. (Round intermediate calculations to 4 decimal places, e.g. 15.2515 and final answers to 0 decimal places, e.g. 2,520.)

Amount
Net income $

Entry field with incorrect answer

Break-even point $

Entry field with incorrect answer
Answer #1
Current Proposed-1 Proposed-2
Selling price per unit 5        1,771,875        2,052,000
Variable costs
Cost of goods sold per unit 2.5            928,125        1,080,000
Selling Expenses per unit 0.49            165,375            254,880
Admin Expense per unit 0.26              87,750            112,320
Contribution Margin per unit 1.75            590,625            604,800
Fixed costs
COGS 225000            225,000            225,000
Selling Expense 182700            182,700            225,700
Admin Expense 105300            105,300            105,300
Net Operating Income              77,625              48,800
CM Ratio = CM/Sales 33.3333% 29.4737%
Break even point = Fixed costs/CM Ratio        1,539,000        1,886,429

- Current =1350000/270000 Proposed-1 =5.25*270000*1.25 Proposed-2 =(B2-0.25)*270000*1.6 =900000*75%/270000 =(B4+0.25)*270000*